A ladder, whish measures 14 m positioned against the wall of a house. If the top of the ladder is 6 m above the base of the wall, how far is the bottom of the ladder placed from the base of the wall?
Answer by nerdybill(7384)

A ladder, whish measures 14 m positioned against the wall of a house. If the top of the ladder is 6 m above the base of the wall, how far is the bottom of the ladder placed from the base of the wall?
.
The ladder is the hypotenuse.
.
Let x = distance from wall
then
x^2 + 6^2 = 14^2
x^2 + 36 = 196
x^2 = 160
x = 12.65 m